OpenStudy (anonymous):

What is the derivative of e^(sin2x)? Is it 2cos(2x)e^(sin2x)?

zepdrix (zepdrix):

Yes c: Oh closed... figure it out already?

OpenStudy (anonymous):

Actually, no, quick question! Where did the 2 in front of the cos come from? @zepdrix

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\left[e^{\sin(2x)}\right]'}\]So we have to apply the chain rule a bunch of times here :d

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\left[e^{\sin(2x)}\right]'}=e^{\sin(2x)}\color{royalblue}{\left[\sin(2x)\right]'}\]So there is our first chain being applied, we need the derivative of this inner function sin(2x).

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{\left[e^{\sin(2x)}\right]'}=e^{\sin(2x)}\cos(2x)\color{royalblue}{\left[2x\right]'}\]Then chain one last time :)

zepdrix (zepdrix):

See where that pesky 2 is coming from?

zepdrix (zepdrix):

Hopefully the coloring made sense. Blue represents something we need to differentiate*

OpenStudy (anonymous):

So is the [2x]' from the (2x) in the exponent?

zepdrix (zepdrix):

I guess it's more accurate to say that it's coming from the sin(2x). The (2x) inside of the sine function let's us know we have to chain rule again.

OpenStudy (anonymous):

Got it! Thanks! :)

zepdrix (zepdrix):

It's coming from this guy: \(\Large\rm \color{royalblue}{\left[\sin(2x)\right]'}\) After differentiating the outer function (sine), we have to chain.

zepdrix (zepdrix):

Chain rule is really freaky and weird! Once you get a good handle on it though, you'll love it. It'll be your best friend.

OpenStudy (anonymous):

I'll try to make amends with it. :)